Divide and Simplify
10a^9 then divide by 9a^4
2 then divide by 6
10a^9 divide 9a^4
----- divide-----
2 divide 6
I sure hope you understand what I am trying to say.

Answer by london maths tutor(243)

(10a^9/9a^4) ÷ 2/6
= (10/9)a^5 * (6/2)
= (10/9)a^5 * 3
= (30/9)a^5
= (3 3/9)a^5
= (1 1/3)a^5